What is the equation of the line passing through the point (2,1) and perpendicular to the line with the equation 3x-6y=1?
Answer by Alan3354(69443)

What is the equation of the line passing through the point (2,1) and perpendicular to the line with the equation 3x-6y=1?
---------------
Find the slope m of the line
Put into standard slope intercept form y = mx+b
3x-6y=1
-6y = -3x + 1
y = (1/2)x -1/6
So m = 1/2
A line perpendicular has a negative inverse slope, -1/m
So m = -2
Use the slope-intercept form again
y-y1 = m*(x-x1)
y-1 = -2*(x-2)
y-1 = -2x+4
2x+y = 5